# Aouad, Ji, and Shaposhnik

### v1

@@ 4 Simple Approximately Optimal Policies
We establish that computing optimal policies for PSI is NP-hard (Section 4.1). We therefore turn our attention to the development and analysis of simple approximate solutions that utilize relaxations that approximate the value function. In Section 4.2, we propose two approximation methods: Whittle’s integral and the free-info relaxation, which adapt tools from related literature. Interestingly, we show that Whittle’s integral provides a tighter upper bound, and establish the asymptotic optimality of the index policy based on the opening thresholds. In general, we exploit these problem relaxations in Section 4.3 to devise approximation algorithms, in particular achieving a $(1-1/e)$-approximation ratio, through a reduction to the maximization of stochastic submodular functions (Asadpour and Nazerzadeh, 2015, Singla, 2018, Beyhaghi and Kleinberg, 2019).
@@ 4.2 Relaxations and upper bounds
The free-info relaxation considers an alternative setting, in which there is no cost of P-opening and F-opening, but instead, the rewards of the boxes are capped according to the threshold values. The notion of a feasible policy does not change in this new formulation, only the cost accounting is modified. Specifically, for every box $i\in\mathcal{C}$, we define the capped value of F-opening to be $K_{i}=\min\{V_{i},\sigma^{F}_{i}\}$ and the capped value of P-opening to be $\widetilde{K}_{i}=\min\{V_{i},\sigma^{F\mid T_{i}}_{i},\sigma^{P}_{i}\}$. Similarly, for every box $i\in{\cal P}$, we define the capped value of type $t_{i}$ as $K_{i}^{t_{i}}=\min\{\sigma_{i}^{F\mid t_{i}},V_{i}\}$. Consequently, in the free-info formulation, the expected reward of policy $\pi$ in state $({\cal C},{\cal P},y)$ is
@@
Recall that we interpret the caps as fair strike prices of call options on the box values, whose expected payoff offsets the opening costs. The formulation (17) can be viewed as an alternative cost accounting where Pandora opens boxes free of charge (there are no negative terms), and in exchange, there is a short position on a call option that exactly compensates the opening costs. However, Pandora has the advantage of being subjected to the short contract only if the box is ultimately selected; otherwise, the short contract is void. Under this optimistic assumption, Pandora essentially collects one of the capped rewards, $K_{i}$, $\widetilde{K}_{i}$, or $\widetilde{K}^{t_{i}}_{i}$, depending on the chosen opening mode.
Going one step further, the free-info relaxation in state $(\mathcal{C},\mathcal{P},y)$ corresponds to the following quantity:
$J^{K}(\mathcal{C},\mathcal{P},y)=\mathbb{E}\left[\max\left\{y,\underset{i\in\mathcal{C}}{\max}\left\{\max\left\{K_{i},\widetilde{K}_{i}\right\}\right\},\underset{i\in\mathcal{P}}{\max}\left\{K_{i}^{t_{i}}\right\}\right\}\right]\ .$ (18)
Similarly to Whittle’s integral, we eliminate the dependence on policy $\pi$ by choosing the box with the highest capped reward. Therefore, $J^{K}(\mathcal{C},\mathcal{P},y)$ is an upper bound on (17), which might not be attainable through a specific policy $\pi$.
We are now ready to state our main result with respect to Whittle’s integral and the free-info relaxation, showing that both methods provide valid upper bounds on the value function.
@@
The proof appears in Appendix 9.1. These upper bounds have been independently derived in the literature for related optimization settings. To our knowledge, this result is the first to formally connect these methods and show that Whittle’s integral provides a tighter upper bound. The proof relies on results by Brown and Smith, (2013) who exhumed an unnoticed technical claim in Whittle, (1980) to derive Whittle’s integral (for the cost-discounted case), and Beyhaghi and Kleinberg, (2019) who proposed the free-info relaxation; a technical difference in our setting is that information can be acquired sequentially over multiple stages.
We note in passing that Lemma 4.2 can be used to establish the asymptotic optimality of the index policy that prioritizes the highest threshold inspection. For brevity, we relegate this result to Appendix 9.2.
@@ 4.3 Near-optimal algorithms
In this section, we leverage the upper bounds of Section 4.2 to construct simple near-optimal policies. Specifically, we develop an algorithm with a worst-case approximation ratio of $(1-1/e)$ using a reduction to the stochastic submodular optimization problem.
Our approach is an application of the framework by Singla, (2018) and Beyhaghi and Kleinberg, (2019). In this context, the main idea is to focus on policies for which the free-info upper bound $J^{\pi,K}$ of equation (17) is exactly equal to the value function $J^{\pi}$ of equation (16). It turns out there is a simple recipe for constructing such policies: (i) commit before opening any box on the selection mode of each box (i.e., whether we start by F-opening or P-opening it), and (ii) strictly follow the order of decreasing opening thresholds. That is, we partition the set of boxes into two subsets, $[N]=F\cup P$, such that boxes $i\in F$ are initially $F$-opened, whereas boxes $i\in P$ are $P$-opened, if inspected. Next, in each state, we select the action with the highest associated threshold. The resulting class of policies $\pi^{F,P}$ is termed committing policies. The reader is referred to the Algorithm 1 for a complete description.
@@
Intuitively, committing policies are not exposed to any cost accounting error, in comparing the opening costs to the corresponding short positions on call options. Recall that this free-info cost accounting is optimistic because Pandora is subjected to the short contract only if the box is ultimately selected. In the proof of Lemma 4.3, we identify necessary and sufficient conditions such that Pandora always selects any F-opened box (or F-opens any P-opened box) whenever the call option (or two-stage call option) is in the money. This way, Pandora meets the contractual payment of the option, which is equal to the opening costs in expectation. In particular, we show these necessary and sufficient conditions are met by committing policies.
Now, this introduction of such committing policies raises the following question: how should we design the partition $(F,P)$? Can we construct committing policies that achieve strong (constant-factor) approximation ratios? Based on the expected profit characterization of Lemma 4.3, we can compare committing policy $\pi^{F,P}$’s payoff $\mathbb{E}[\max\{\underset{i\in F}{\max}\{K_{i}\},\underset{i\in P}{\max}\{\widetilde{K}_{i}\}\}]$ to the free-info upper-bound $J^{K}(s_{0})=\mathbb{E}[\underset{i\in[N]}{\max}\max\{K_{i},\widetilde{K}_{i}\}]$ (see Equation (18)). The former ex-ante chooses either $K_{i}$ or $\widetilde{K}_{i}$, before observing their realizations, depending on whether $i\in F$ or $i\in P$. By contrast, the latter chooses the maximum reward $\max\left\{K_{i},\widetilde{K}_{i}\right\}$ out of the two opening modes after seeing their realizations. This informal comparison leads to the design of simple approximation algorithms.
First, note that the following upper bound on the free-info setting always holds:
@@
where the inequality holds since $\max\{a,b\}\leq\max\{c,d\}+\max\{e,f\}$ for every scalars $a\leq\max\{c,e\}$ and $b\leq\max\{d,f\}$, and the last equality proceeds from Lemma 4.3. Importantly, the above inequality implies that, for any partition of $(F,P)$ of $[N]$, either the committing policy $\pi^{F,P}$ or the one $\pi^{P,F}$ that flips the opening modes achieves a factor-$\frac{1}{2}$ approximation of the optimum.
More broadly, the performance of a well-designed committing policy is related to the adaptivity gap in stochastic submodular maximization; the question can be reformulated as optimizing a certain submodular function over random entries and the goal is to bound the gap between non-adaptive policies, that select those entries before their realizations are revealed, and adaptive ones. By leveraging fundamental results in this area, we obtain the following theorem.
@@
In Appendix 9.5, we devise a reduction from Pandora’s box problem to the problem of maximizing a stochastic submodular function under additional constraints. In this context, Asadpour and Nazerzadeh, (2015) showed, among other results, that the adaptivity gap (i.e., the performance gap between non-adaptive and adaptive policies) is lower bounded by $1-1/e$, and that an approximate non-adaptive policy can be efficiently constructed using a continuous-greedy algorithm.
@@ 5 Numerical Study
We conduct numerical experiments to further explore the characteristics of optimal policies and assess the effectiveness of natural candidate policies. Section 5.1 describes the process for generating a diverse set of problem instances. Section 5.2 quantifies the extent to which our theoretical characterization captures optimal decisions. Section 5.3 compares the performance of various policies inspired by theoretical analysis. In Section 5.4, we examine factors that increase the tendency of the optimal policies to P-open. Finally, in Section 5.5 we measure the capacity of index policies for representing optimal decisions and propose a more general class of policies.
@@ 5.1 Problem instances
Our design of numerical experiments is based on the observation that the interaction between opening thresholds significantly impacts the behavior of optimal policies. We therefore create instances that consist of boxes with diverse values of opening thresholds.
We begin by generating a collection of “prototypical” boxes. To create a prototypical box $i$, we sample 5 values for the support of $V_{i}$ from the uniform distribution $\mathcal{U}(0,10)$; every type $T_{i}$ can realize up to at most three types (that is, $\mid\Gamma_{i}\mid\leq 3$); for each type we uniformly sample the conditional distribution $V_{i}\mid T_{i}=t$ over the support; we sample the F-opening and P-opening costs from the uniform distributions $\mathcal{U}(0,5)$ and $\mathcal{U}(0,3)$, respectively. We then generate a candidate set of prototypical boxes out of which we select a small representative set (Figure 3).
Each problem instance consists of randomized copies of the prototypical boxes, in varying compositions with $N=2$ to $N=16$ boxes. Due to running time limitations, there are 1000 instances whose size is 2 to 9 (“small instances” which can be solved to optimality using DP), and there are 300 instances whose size is 10 to 16 (“large instances” which can only be solved using heuristics). To avoid trivial cases, we only generate boxes where $c^{F}_{i}\geq c^{P}_{i}$ (which is why in Figure 3 most prototypical boxes have greater P-thresholds).
@@ 5.2 Quantifying the coverage of analytical results
The analysis of Section 3 revealed various insights about the nature of the optimal policy. Since the analysis includes sufficient but not always necessary conditions, we are interested in understanding to what extent the sufficient conditions of Theorems 1–3 describe an optimal policy.
To this end, we conduct two types of experiments. First, we examine optimal solutions to the “small instances” and compute the percentage of states where the sufficient conditions of Theorems 1–3 hold (i.e., where our theorems prescribe optimal actions). We implement the DP recursion in (5) and compute coverage rates, defined as the percentages of system states where the optimal action can be determined by our sufficient conditions.
The results are summarized in Table 1. The overall coverage rate is presented in the 5th column, and is divided into specific conditions in the 2nd to 4th columns. Looking at the 5th column of the table, we see that as the problem increases in size, the sufficient conditions of Theorems 1–3 capture optimal actions in all but a relatively few states. The last two columns present the recall of Theorems 2 and 3. These reflect the percentage of states in which the sufficient conditions of the theorems hold out of all states where F-opening (column 6) and P-opening (column 7) are optimal. For example, for instances of size 5 (4th row in the table), Theorem 2 captures 97% of states where F-opening is optimal, and Theorem 3 captures 77% of all states where P-opening is optimal. We observe that identifying P-opening is more challenging than identifying when F-opening is optimal, but overall, the experiment suggests that our characterization explains to a large degree the behavior of the optimal policy. Note that the recall of Theorem 1 is 100% as the theorem perfectly describes when stopping is optimal.
We conduct a second experiment where we compare two DP implementations that recursively compute the optimal policy. The first policy “Naïve DP” is implemented using the DP recursion in Equation (5), which does not leverage any structure of the optimal policy. The second implementation, on the other hand, utilizes the sufficient conditions of Theorems 1–3 which determine optimal actions and do not require calling the DP recursion when the conditions are satisfied. Table 2 compares the two approaches in terms of the total number of states created, the total number of visits to system states, and the runtime. We see that the characterization of the optimal policy not only provides insights about the optimal policy, but also bears computational benefits, allowing to solve problems at significantly shorter times. That said, the “curse of dimensionality” cannot be avoided, as can be observed by the rate at which the size of the state space grows.
@@ 5.3 Policy benchmark and computational study
  • $\pi^{\textrm{W}}$ – a one-step lookahead policy (Bertsekas, 2019) based on approximate value functions using Whittle’s integral. Another natural policy to consider is the one-step lookahead policy that uses $J^{K}$ to approximate the value function. However, given that $J^{W}$ provides a tighter upper bound (Lemma 4.2), we only test $\pi^{W}$ in the experiments.
  • $\pi^{STP}$ – a policy that determines its next action based on the maximal myopic improvement (using a single “test,” hence Single Test Policy) in comparison to stopping. Computing the improvement from each non-open box becomes tractable by ignoring other non-open boxes, and follows the optimal policy for the single box problem (see Figure 2).
To evaluate performance, we measure the value function and runtime of each problem instance under different policies, starting from the initial state. As the expected profit can vary significantly among different instances, we normalize the expected profit of each policy relative to the optimal value obtained by policy $\pi^{\textrm{OPT}}$. For large instances where the optimal policy cannot be computed, we use the best-performing policy to normalize the expected profit. Thus, the performance of each policy is reported as a number between 0 and 1, where larger values indicate better performance. For each policy and a set of instances, we report on performance using the average across instances and the worst instance tested.
Average performance. All heuristics perform quite well in terms of their average performance, often exceeding 90% of OPT. However, $\pi^{\textrm{index}}$ and $\pi^{F^{*},P^{*}}$ are superior, with average performances that are close to 100%.
Worst-case performance. We observe that $\pi^{\textrm{W}}$ can perform poorly, at times nearing zero profit. This suggests that Whittle’s integral approach may not be a suitable approach to deriving tight bounds for our problem, and perhaps to other Pandora’s box problems as well. We also see that $\pi^{\textrm{index}}$ is superior to $\pi^{STP}$ and that its relative performance improves as the number of boxes increases. We observe that $\pi^{\textrm{index}}$ approaches optimality as $N$ increases, which is aligned with its asymptotic optimality. Outperforming all others is policy $\pi^{F^{*},P^{*}}$, which achieves the best worst-case performance, with a loss of at most 7% for instances of size 2. This suggests that index policies based on our opening thresholds are effective for solving the problem. We note, however, that this policy is not optimal, and finding the optimal committing policy is computationally challenging (an approximation algorithm is provided in Asadpour and Nazerzadeh, 2015).
Runtime. Table 5 in Appendix 10 summarizes the runtimes. As expected, the time for solving instances to optimality grows exponentially fast, making it challenging to solve instances with more than 9 boxes. We also see that applying policy $\pi^{\textrm{W}}$ becomes more challenging as the number of boxes increases, due to the calculation of Whittle’s integral. Similarly, the exhaustive enumeration carried out by policy $\pi^{F^{*},P^{*}}$ could result in long runtimes. We note that the runtime also depends on the size of the support of the distribution (which we assumed to be 5) and the number of box types (which we assumed to be at most 3). Note that these runtimes include the entire calculation of the value function and that simply applying these policies is very fast.
Optimality. Table 4 shows the percentage of instances for which different policies attain the optimal expected objective value. In contrast to the near-optimal performance displayed in Table 3, policy $\pi^{\textrm{W}}$ is almost never optimal. Additionally, even policies $\pi^{F,P}$ and $\pi^{\textrm{index}}$, which perform better, in many cases do not achieve the optimal objective value. We notice that overall, the performance of the latter two index policies improves with the number of boxes, which aligns with the asymptotic optimality result (Corollary 9.1).

### v2

@@ 4 Simple Approximately Optimal Policies
We establish that computing optimal policies for PSI is NP-hard (Section 4.1). We therefore turn our attention to the development and analysis of simple approximate solutions that utilize relaxations that approximate the value function. In Section 4.2, we propose two approximation methods: Whittle’s integral and the free-info relaxation, which adapt tools from related literature. Interestingly, we show that Whittle’s integral provides a tighter upper bound, and establish the asymptotic optimality of the index policy based on the opening thresholds. In general, we exploit these problem relaxations in Section 4.3 to devise approximation algorithms, in particular achieving a (11/e)(1-1/e)-approximation ratio, through a reduction to the maximization of stochastic submodular functions (Asadpour and Nazerzadeh, 2015, Singla, 2018, Beyhaghi and Kleinberg, 2019).
@@ 4.1 Hardness of the problem
Finally, our reduction from subset partition has some similarities with the NP-hardness proof of Fu et al., (2023) in a related computational setting; see Remark 10.13 in Appendix 10.6 where we discuss similarities and important differences in both reduction proofs.
@@ 4.2 Relaxations and upper bounds
The free-info relaxation considers an alternative setting, in which there is no cost of P-opening and F-opening, but instead, the rewards of the boxes are capped according to the threshold values. The notion of a feasible policy does not change in this new formulation, only the cost accounting is modified. Specifically, for every box $i\in\mathcal{C}$, we define the capped value of F-opening to be $K_{i}=\min\{V_{i},\sigma^{F}_{i}\}$ and the capped value of P-opening to be $\widetilde{K}_{i}=\min\{V_{i},\sigma^{F\mid T_{i}}_{i},\sigma^{P}_{i}\}$. Similarly, for every box $i\in{\cal P}$, we define the capped value of type $t_{i}$ as $\widetilde{K}_{i}^{t_{i}}=\min\{\sigma_{i}^{F\mid t_{i}},V_{i}\}$. Consequently, in the free-info formulation, the expected reward of policy $\pi$ in state $({\cal C},{\cal P},y)$ is
@@
Recall that we interpret the caps as fair strike prices of call options on the box values, whose expected payoff offsets the opening costs. The formulation (18) can be viewed as an alternative cost accounting where Pandora opens boxes free of charge (there are no negative terms), and in exchange, there is a short position on a call option that exactly compensates the opening costs. However, Pandora has the advantage of being subjected to the short contract only if the box is ultimately selected; otherwise, the short contract is void. Under this optimistic assumption, Pandora essentially collects one of the capped rewards, $K_{i}$, $\widetilde{K}_{i}$, or $\widetilde{K}^{t_{i}}_{i}$, depending on the chosen opening mode. Going one step further, the free-info relaxation in state $(\mathcal{C},\mathcal{P},y)$ corresponds to the following quantity:
$J^{K}(\mathcal{C},\mathcal{P},y)=\mathbb{E}\left[\max\left\{y,\underset{i\in\mathcal{C}}{\max}\left\{\max\left\{K_{i},\widetilde{K}_{i}\right\}\right\},\underset{i\in\mathcal{P}}{\max}\left\{\widetilde{K}_{i}^{t_{i}}\right\}\right\}\right]\ .$ (19)
Similarly to Whittle’s integral, we eliminate the dependence on policy $\pi$ by choosing the box with the highest capped reward. Therefore, $J^{K}(\mathcal{C},\mathcal{P},y)$ is an upper bound on (18), which might not be attainable through a specific policy $\pi$. We are now ready to state our main result with respect to Whittle’s integral and the free-info relaxation, showing that both methods provide valid upper bounds on the value function.
@@
The proof appears in Appendix 10.1. These upper bounds have been independently derived in the literature for related optimization settings. To our knowledge, this result is the first to formally connect these methods and show that Whittle’s integral provides a tighter upper bound. The proof relies on results by Brown and Smith, (2013) who exhumed an unnoticed technical claim in Whittle, (1980) to derive Whittle’s integral (for the cost-discounted case), and Beyhaghi and Kleinberg, (2019) who proposed the free-info relaxation; a technical difference in our setting is that information can be acquired sequentially over multiple stages. We note in passing that Lemma 4.2 can be used to establish the asymptotic optimality of the index policy that prioritizes the highest threshold inspection. For brevity, we relegate this result to Appendix 10.2.
@@ 4.3 Near-optimal algorithms
In this section, we leverage the upper bounds of Section 4.2 to construct simple near-optimal policies. Specifically, we develop an algorithm with a worst-case approximation ratio of (11/e)(1-1/e) using a reduction to the stochastic submodular optimization problem. Our approach is an application of the framework by Singla, (2018) and Beyhaghi and Kleinberg, (2019). In this context, the main idea is to focus on policies for which the free-info upper bound $J^{\pi,K}$ of equation (18) is exactly equal to the value function $J^{\pi}$ of equation (17). It turns out there is a simple recipe for constructing such policies: (i) commit before opening any box on the selection mode of each box (i.e., whether we start by F-opening or P-opening it), and (ii) strictly follow the order of decreasing opening thresholds. That is, we partition the set of boxes into two subsets, $N]=F\cup P$, such that boxes $i\in F$ are initially $F$-opened, whereas boxes $i\in P$ are $P$-opened, if inspected. Next, in each state, we select the action with the highest associated threshold. The resulting class of policies $\pi^{F,P}$ is termed committing policies. The reader is referred to the Algorithm 1 for a complete description.
@@
Intuitively, committing policies are not exposed to any cost accounting error, in comparing the opening costs to the corresponding short positions on call options. Recall that this free-info cost accounting is optimistic because Pandora is subjected to the short contract only if the box is ultimately selected. In the proof of Lemma 4.3, we identify necessary and sufficient conditions such that Pandora always selects any F-opened box (or F-opens any P-opened box) whenever the call option (or two-stage call option) is in the money. This way, Pandora meets the contractual payment of the option, which is equal to the opening costs in expectation. In particular, we show these necessary and sufficient conditions are met by committing policies. Now, this introduction of such committing policies raises the following question: how should we design the partition $(F,P)$? Can we construct committing policies that achieve strong (constant-factor) approximation ratios? Based on the expected profit characterization of Lemma 4.3, we can compare committing policy $\pi^{F,P}$’s payoff $\mathbb{E}[\max\{\underset{i\in F}{\max}\{K_{i}\},\underset{i\in P}{\max}\{\widetilde{K}_{i}\}\}]$ to the free-info upper-bound $J^{K}(s_{0})=\mathbb{E}[\underset{i\in[N]}{\max}\max\{K_{i},\widetilde{K}_{i}\}]$ (see Equation (19)). The former ex-ante chooses either $K_{i}$ or $\widetilde{K}_{i}$, before observing their realizations, depending on whether $i\in F$ or $i\in P$. By contrast, the latter chooses the maximum reward $\max\left\{K_{i},\widetilde{K}_{i}\right\}$ out of the two opening modes after seeing their realizations. This informal comparison leads to the design of simple approximation algorithms. First, note that the following upper bound on the free-info setting always holds:
@@
where the inequality holds since $\max\{a,b\}\leq\max\{c,d\}+\max\{e,f\}$ for every scalars $a\leq\max\{c,e\}$ and $b\leq\max\{d,f\}$, and the last equality proceeds from Lemma 4.3. Importantly, the above inequality implies that, for any partition of $(F,P)$ of $[N]$, either the committing policy $\pi^{F,P}$ or the one $\pi^{P,F}$ that flips the opening modes achieves a factor-$\frac{1}{2}$ approximation of the optimum. More broadly, the performance of a well-designed committing policy is related to the adaptivity gap in stochastic submodular maximization; the question can be reformulated as optimizing a certain submodular function over random entries and the goal is to bound the gap between non-adaptive policies, that select those entries before their realizations are revealed, and adaptive ones. By leveraging fundamental results in this area, we obtain the following theorem.
@@
In Appendix 10.5, we devise a reduction from Pandora’s box problem to the problem of maximizing a stochastic submodular function under additional constraints. In this context, Asadpour and Nazerzadeh, (2015) showed, among other results, that the adaptivity gap (i.e., the performance gap between non-adaptive and adaptive policies) is lower bounded by 11/e1-1/e, and that an approximate non-adaptive policy can be efficiently constructed using a continuous-greedy algorithm.
@@ 5 Numerical Study
We conduct numerical experiments to further explore the characteristics of optimal policies and assess the effectiveness of natural candidate policies. Section 5.1 describes the process for generating a diverse set of problem instances. Section 5.2 quantifies the extent to which our theoretical characterization captures optimal decisions. Section 5.3 compares the performance of various policies inspired by theoretical analysis. Finally, in Section 5.4, we examine factors that increase the tendency of the optimal policies to P-open.
@@ 5.1 Problem instances
Our design of numerical experiments is based on the observation that the interaction between opening thresholds significantly impacts the behavior of optimal policies. We therefore create instances that consist of boxes with diverse values of opening thresholds. We begin by generating a collection of “prototypical” boxes. To create a prototypical box $i$, we sample 5 values for the support of $V_{i}$ from the uniform distribution $\mathcal{U}(0,10)$; every type $T_{i}$ can realize up to at most three types (that is, $\mid\Gamma_{i}\mid\leq 3$); for each type we uniformly sample the conditional distribution $V_{i}\mid T_{i}=t$ over the support; we sample the F-opening and P-opening costs from the uniform distributions $\mathcal{U}(0,5)$ and $\mathcal{U}(0,3)$, respectively. We then generate a candidate set of prototypical boxes out of which we select a small representative set (Figure 3). Each problem instance consists of randomized copies of the prototypical boxes, in varying compositions with $N=2$ to $N=16$ boxes. Due to running time limitations, there are 1000 instances whose size is 2 to 9 (“small instances” which can be solved to optimality using DP), and there are 300 instances whose size is 10 to 16 (“large instances” which can only be solved using heuristics). To avoid trivial cases, we only generate boxes where $c^{F}_{i}\geq c^{P}_{i}$ (which is why in Figure 3 most prototypical boxes have greater P-thresholds).
@@ 5.2 Quantifying the coverage of analytical results
The analysis of Section 3 revealed various insights about the nature of the optimal policy. Since the analysis includes sufficient but not always necessary conditions, we are interested in understanding to what extent the sufficient conditions of Theorems 1–3 describe an optimal policy. To this end, we conduct two types of experiments. First, we examine optimal solutions to the “small instances” and compute the percentage of states where the sufficient conditions of Theorems 1–3 hold (i.e., where our theorems prescribe optimal actions). We implement the DP recursion in (5) and compute coverage rates, defined as the percentages of system states where the optimal action can be determined by our sufficient conditions. The results are summarized in Table 1. The overall coverage rate is presented in the 5th column, and is divided into specific conditions in the 2nd to 4th columns. Looking at the 5th column of the table, we see that as the problem increases in size, the sufficient conditions of Theorems 1–3 capture optimal actions in all but a relatively few states. The last two columns present the recall of Theorems 2 and 3. These reflect the percentage of states in which the sufficient conditions of the theorems hold out of all states where F-opening (column 6) and P-opening (column 7) are optimal. For example, for instances of size 5 (4th row in the table), Theorem 2 captures 97% of states where F-opening is optimal, and Theorem 3 captures 77% of all states where P-opening is optimal. We observe that identifying P-opening is more challenging than identifying when F-opening is optimal, but overall, the experiment suggests that our characterization explains to a large degree the behavior of the optimal policy. Note that the recall of Theorem 1 is 100% as the theorem perfectly describes when stopping is optimal. We conduct a second experiment where we compare two DP implementations that recursively compute the optimal policy. The first policy “Naïve DP” is implemented using the DP recursion in Equation (5), which does not leverage any structure of the optimal policy. The second implementation, on the other hand, utilizes the sufficient conditions of Theorems 1–3 which determine optimal actions and do not require calling the DP recursion when the conditions are satisfied. Table 2 compares the two approaches in terms of the total number of states created, the total number of visits to system states, and the runtime. We see that the characterization of the optimal policy not only provides insights about the optimal policy, but also bears computational benefits, allowing to solve problems at significantly shorter times. That said, the “curse of dimensionality” cannot be avoided, as can be observed by the rate at which the size of the state space grows.
@@ 5.3 Policy benchmark and computational study
  • $\pi^{\textrm{W}}$ – a one-step lookahead policy (Bertsekas, 2019) based on approximate value functions using Whittle’s integral.
Another natural policy to consider is the one-step lookahead policy that uses $J^{K}$ to approximate the value function. However, given that $J^{W}$ provides a tighter upper bound (Lemma 4.2), we only test $\pi^{W}$ in the experiments. $\pi^{STP}$ – a policy that determines its next action based on the maximal myopic improvement (using a single “test,” hence Single Test Policy) in comparison to stopping. Computing the improvement from each non-open box becomes tractable by ignoring other non-open boxes, and follows the optimal policy for the single box problem (see Figure 2). To evaluate performance, we measure the value function and runtime of each problem instance under different policies, starting from the initial state. As the expected profit can vary significantly among different instances, we normalize the expected profit of each policy relative to the optimal value obtained by policy $\pi^{\textrm{OPT}}$. For large instances where the optimal policy cannot be computed, we use the best-performing policy to normalize the expected profit. Thus, the performance of each policy is reported as a number between 0 and 1, where larger values indicate better performance. For each policy and a set of instances, we report on performance using the average across instances and the worst instance tested. Average performance. All heuristics perform quite well in terms of their average performance, often exceeding 90% of OPT. However, $\pi^{\textrm{index}}$ and $\pi^{F^{*},P^{*}}$ are superior, with average performances that are close to 100%. Worst-case performance. We observe that $\pi^{\textrm{W}}$ can perform poorly, at times nearing zero profit. This suggests that Whittle’s integral approach may not be a suitable approach to deriving tight bounds for our problem, and perhaps to other Pandora’s box problems as well. We also see that $\pi^{\textrm{index}}$ is superior to $\pi^{STP}$ and that its relative performance improves as the number of boxes increases. We observe that $\pi^{\textrm{index}}$ approaches optimality as $N$ increases, which is aligned with its asymptotic optimality. Outperforming all others is policy $\pi^{F^{*},P^{*}}$, which achieves the best worst-case performance, with a loss of at most 7% for instances of size 2. This suggests that index policies based on our opening thresholds are effective for solving the problem. We note, however, that this policy is not optimal, and finding the optimal committing policy is computationally challenging (an approximation algorithm is provided in Asadpour and Nazerzadeh, 2015). Runtime. Table 5 in Appendix 11 summarizes the runtimes. As expected, the time for solving instances to optimality grows exponentially fast, making it challenging to solve instances with more than 9 boxes. We also see that applying policy $\pi^{\textrm{W}}$ becomes more challenging as the number of boxes increases, due to the calculation of Whittle’s integral. Similarly, the exhaustive enumeration carried out by policy $\pi^{F^{*},P^{*}}$ could result in long runtimes. We note that the runtime also depends on the size of the support of the distribution (which we assumed to be 5) and the number of box types (which we assumed to be at most 3). Note that these runtimes include the entire calculation of the value function and that simply applying these policies is very fast. Optimality. Table 4 shows the percentage of instances for which different policies attain the optimal expected objective value. In contrast to the near-optimal performance displayed in Table 3, policy $\pi^{\textrm{W}}$ is almost never optimal. Additionally, even policies $\pi^{F,P}$ and $\pi^{\textrm{index}}$, which perform better, in many cases do not achieve the optimal objective value. We notice that overall, the performance of the latter two index policies improves with the number of boxes, which aligns with the asymptotic optimality result (Corollary 10.1).